MOA-II Obsevation Log Monthly Summary 2008/02
Return to index
Previous month
Next month
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | 11438-11572 | 135 | 726 | 43560 |
LMC | 2412-2523 | 112 | 606 | 177300 |
SMC | 597-609 | 13 | 26 | 7680 |
ALERT | - | 0 | 0 | 0 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 260 | 1358 | 228540 (63h29m00s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
51 | 50 | 47 | 37 | 114 | 23 | 34 | 39 | 108 | 21 | 24 | 21 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
20 | 17 | 17 | 17 | 17 | 17 | 15 | 15 | 13 | 9 | 0 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
23 | 22 | 22 | 21 | 21 | 25 | 112 | 110 | 107 | 33 | 25 | 25 |
Field13 | Field14 | Field15 | Field16 |
29 | 31 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
13 | 13 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
SPECIAL TARGET NAME
CCD TEMPERATURE
Previous month
Next month

VACUUM
Previous month
Next month

TELESCOPE TEMPERATURE
Previous month
Next month

COMMENTS
2008/02/1: No observation. Pump off. / night time = 24720(sec)
2008/02/2: DARK D3394(300sec), D3395(60sec) are OK. / night time = 25020(sec)
2008/02/3: DARK D3401(300sec) is OK. / night time = 25260(sec)
2008/02/4: DARK D3411(300sec) is OK. / night time = 25500(sec)
2008/02/5: DARK D3414(300sec), D3416(60sec) are OK. / night time = 25740(sec)
2008/02/6: DARK D3419(300sec), D3420(60sec) are OK. / night time = 25980(sec)
2008/02/7: DARK D3429(300sec), D3430(60sec) are OK. / night time = 26220(sec)
2008/02/8: DARK D3432(60sec) is OK. / night time = 26460(sec)
2008/02/9: No observation. / night time = 26760(sec)
2008/02/10: I took FLAT(F2168-F2203). DARK D3435(15sec) is OK. / night time = 27000(sec)
2008/02/11: No observation. / night time = 27240(sec)
2008/02/12: DARK D3438(300sec), D3439(60sec) are OK. / night time = 27480(sec)
2008/02/13: No observation. / night time = 27720(sec)
2008/02/14: No observation. / night time = 27960(sec)
2008/02/15: No observation. / night time = 28200(sec)
2008/02/16: DARK D3443(300sec), D3445(60sec), D3446(600sec:for the record) are OK. I took the special run using only V-filter because it was bad seeing. / night time = 28440(sec)
2008/02/17: DARK D3452(600sec), D3453(300sec), D3454(60sec) are OK. I took the stream(V,I). / night time = 28680(sec)
2008/02/18: DARK D3457(300sec), D3459(60sec) are OK. / night time = 28920(sec)
2008/02/19: DARK D3465(240sec), D3466(60sec) are OK. I changed exposure time(240sec) due to MOON. / night time = 29100(sec)
2008/02/20: DARK D3473(240sec), D3474(60sec) are OK. I changed exposure time(240sec) due to MOON. / night time = 29400(sec)
2008/02/21: No observation. / night time = 29640(sec)
2008/02/22: No observation. We work on building raid on moa02 and moatel. / night time = 29880(sec)
2008/02/23: No observation. We work on building raid on moa02 and moatel. / night time = 30120(sec)
2008/02/24: No observation. We have completed building raid on moa02 and moatel. / night time = 30360(sec)
2008/02/25: telescope controler problem after 4am / night time = 30600(sec)
2008/02/26: D3481-60, D3482-300 are ok. telescopeM1 stopped again at 15:02 UTC / night time = 30780(sec)
2008/02/27: D3483-300 is ok / night time = 31020(sec)
2008/02/28: No Obs / night time = 31260(sec)
2008/02/29: No Obs. / night time = 31500(sec)
TOTAL NIGHT TIME: 816960(sec)
OPERATING RATE: 34.6 %
GB: 8.9 %
LMC: 24.7 %
SMC: 1.1 %
ALERT: 0.0 %
GRB: 0.0 %
Return to index
Previous month
Next month